OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14 No solutions x = -1, x = 8.3 x = 0, x = 7 x = -1, x = 7

OpenStudy (anonymous):

please help me <3

OpenStudy (superdavesuper):

subtract 2 then divide by 3 on both the left n right hand side of the eqn....what will u get?

OpenStudy (anonymous):

u can't divide 14 by 3 cuz it'll be a messed up #

OpenStudy (superdavesuper):

subtract 2 first plz....

OpenStudy (anonymous):

ohhh!!! 12

OpenStudy (superdavesuper):

then divide by 3....what would u get?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

4

OpenStudy (superdavesuper):

7 is one of the right answer :)

OpenStudy (anonymous):

and -1 is the other one.

OpenStudy (anonymous):

thxxx

OpenStudy (superdavesuper):

the other is a lil more tricky....do u know what the | | symbols mean?

OpenStudy (superdavesuper):

Oh nevermind lol u are a smart lady :D

OpenStudy (anonymous):

We know that |xI = +/-x So |x - 3| = +/-(x - 3) So first take it (x - 3) that is our expression become 3*(x - 3) + 2 = 14 & solve for x then Use -(x - 3) so expression become 3*(-(x - 3)) + 2 = 14 & solve for x Hope this will help you ..